# thanks: current affiliation: IBM Quantum

### v1

@@ II.2 Illustration of the QEC-LDD Theorem: the $[[4,2,2]]$ code
To illustrate the QEC-LDD construction, we analyze the example of the $[[4,2,2]]$ code (see Appendix A for a review of this code). The stabilizer group of this code is $\mathcal{S}=\langle XXXX,ZZZZ\rangle$. A set of logical operators for the code can be chosen such that $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, i.e., $\mathcal{L}=\langle i,XIIX,IIXX,IIZZ,ZIIZ\rangle$. Up to a global phase, $\{XIIX,IIXX,IIZZ,ZIIZ\}$ is the generator set of the $16$-element decoupling group $\tilde{\mathcal{G}}=\{g_{j}\}$. This LDD group suppresses every logical error $E$ in the $4\times 16=64$-element $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$: the group-averaged logical system-bath Hamiltonian $\langle H_{SB}^{\mathcal{L}}\rangle_{\mathcal{G}}$ vanishes, as every one of its terms is a logical operator that anticommutes with an element of $\mathcal{G}$.
The errors that are not suppressed are easily seen to be the $16$-element set =$\mathcal{E}=\{P\otimes Q\otimes P\otimes P:P,Q\in\{I,X,Y,Z\}\}$, since they commute with all elements of $\mathcal{G}$. Each such error is detectable because it anticommutes with at least one element of $\mathcal{S}$. Moreover, every such error is correctable since it is equal to the product of the stabilizer $P\otimes P\otimes P\otimes P$ and $I\otimes R\otimes I\otimes I$, where $R=Q^{\dagger}P$ can be determined (up to the global phase) from the value of the syndrome. The recovery operation is then the weight-one operator $I\otimes R\otimes I\otimes I$.
Ignoring phases, the Pauli group on $4$ physical qubits, $\tilde{\mathcal{P}}_{4}$, has $2564^{4}=$ 256 elements. Of these, $64$ are the elements of $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$, $16$ are the unsuppressed errors $\mathcal{E}$, which leaves $176$ other elements $[P]=\{\pm P,\pm iP\}\in\tilde{\mathcal{P}}_{4}$ such that corresponding $P\in\mathcal{P}_{4}$ anticommutes with at least one element of $\mathcal{G}$ and is, therefore, suppressed by it.
Appendix C illustrates the unencoding of the $[[4,2,2]]$ code. In this case, $\mathcal{P}_{2}$ decouples all errors on the first two (data) qubits and only $X_{3}$ or $X_{4}$ are left as undecoupled bit-flip errors on the last two qubits.
II.3 Avoiding the use of virtual-$Z$ gates
One of the limitations of our experimental setup is that physical $Z$ gates are unavailable: fixed-frequency transmons utilize virtual-$Z$ gates [94, 95]. Such gates are problematic for DD [95]. We now demonstrate that there are DD sequences satisfying the conditions of the QEC-LDD theorem which do not use any physical $Z$ pulses. To construct such a sequence, we first choose generators of $\mathcal{L}$ that do not involve physical $Z$ gates. An example choice would be $\mathcal{L}=\langle i,XIIX,IIXX,IIYY,YIIY\rangle$. One could follow the construction below with this choice. However, in order to improve the ability of the DD sequence to suppress $ZZ$ crosstalk we suggest a different choice. First, multiply the original generators of $\mathcal{L}$ (that is, $XIIX,IIXX,IIZZ,ZIIZ$) by the stabilizers $XXXX$, $XXXX$, $XXXX$, and $ZZZZ$, respectively, to obtain $\mathcal{L}^{\prime}=\langle i,IXXI,XXII,XXYY,IZZI\rangle$. Then, after some group operations, we can write the same group using a different set of generators containing no $Z$’s: $\mathcal{L}^{\prime}=\langle i,XXII,IIYY,IYIY,XIXI\rangle$. The DD sequence consisting of two repetitions of $XIXI$, $IYIY$, $XIXI$, $IIYY$, $XIXI$, $IYIY$, $XIXI$, $XXII$ then corresponds to the decoupling group /{±$\mathcal{L}^{\prime}/\{\pm 1,\pm i\}$, and involves no $Z$ gates.
This sequence was obtained as follows: denote the generators of $\mathcal{L}$ (or $\mathcal{L}^{\prime}$) as $h_{1},h_{2},h_{3},h_{4}$ and choose any Gray code — a sequence of bitstrings $a_{0},\dots,a_{15}$ such that neighboring bitstrings differ in only a single digit. Finally, pick $g_{j}=h^{a_{j}}$ (for $j=0,\dots,15$), where $a_{j}$ is interpreted as a multiindex (i.e. $g_{j}=\prod_{k=1}^{4}h_{k}^{(a_{j})_{k}}$). The pulses of the DD sequence are $g_{j}g_{j+1}$ for $j=0,\dots,15$.
@@ II.3 Avoiding the use of virtual-$Z$ gates
However, in our experiments we found that a decoupling group generated by a subgroup of $\mathcal{L}$ and a subgroup of $\mathcal{S}$ works better in practice because of its robustness and ability to suppress crosstalk. The next section is dedicated to demonstrating experimentally how the hybrid QEC-LDD protocol results in a significant fidelity enhancement of entangled logical qubits.
@@
III Experimental design and methodology
The ibm_kyiv processor consists of coupled, fixed-frequency transmons [97]. Such qubits exhibit an always-on interaction between adjacent pairs [98]. This crosstalk gives rise to weight-2 error terms that correspond to logical errors with our $[[4,2,2]]$ code choice. Other errors of weight $\geq 2$, if present, would likewise correspond to logical errors. To demonstrate that LDD can suppress all logical errors, we design an experiment that enables the unequivocal detection of such errors. To this end, we use the two logical qubits of the $[[4,2,2]]$ code to create logical Bell states |$\ket{\Phi_{\pm}}=(\ket{\overline{00}}\pm\ket{\overline{11}})/\sqrt{2}$ and |$\ket{\Psi_{\pm}}=(\ket{\overline{01}}\pm\ket{\overline{10}})/\sqrt{2}$. The corresponding encoding circuits $U_{\text{enc},\chi}$, where $\chi\in\{\ket{\Phi_{\pm}},\ket{\Psi_{\pm}}\}$, create two copies of physical Bell states, followed by the application of a logical controlled-NOT ($\overline{\text{CNOT}}$) to generate the logical Bell states (see Appendix A).
To estimate the fidelity of a logical Bell state $\chi$, we start from the physical ground state |$\ket{0000}$, encode into the state of interest by applying $U_{\text{enc},\chi}$, and then let it evolve for some time $\tau$, either freely or subject to DD.
If after time $\tau$ we unencode by applying $U_{\text{enc},\chi}^{\dagger}$, then measuring the bitstring $0000$ would indicate that no error occurred.
Figure 2: Circuit schematic for encoding |$\ket{\Phi_{+}}$ and unencoding |$\ket{\Phi_{-}}$ with physical and logical DD. Physical XY4 dynamical decoupling sequences (yellow boxes) are inserted into the idle gaps of the circuit. The code subspace is protected with LDD sequences. This is the DD-protected version of the circuit shown in Fig. 1(a).
Figure 3: Performance of LDD (dataset 1). The experiments encode the logical Bell states |$\ket{\Phi_{+}}$ in (a) and (b), and |$\ket{\Psi_{+}}$ in (c) and (d). We then unencode in the logical Bell state related through $\overline{Z}_{i}$ on either of the logical qubits; i.e., |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ in (a) and (b), and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ in (c) and (d). In this setting, the frequency of $0110$ outcomes [(a) and (c)] is a measure of the fidelity of the originally encoded state, while $0000$ corresponds to the detection of a $\overline{Z}_{i}$ error [(b) and (d)]. Without any DD, the fidelity is low (green in (a) and (c)) and the probability of logical errors is high [green in (b) and (d)]. Standard LDD sequences (LXX and LXY4; dashed light red and blue) improve the fidelity and logical error probability at short times ($\lesssim 15\mu$s) but their performance declines at longer times: they exhibit coherent errors as seen in the oscillations of the corresponding curves. Robust sequences (RLXX and RLXY4; solid dark red and blue) are the top performers; their ability to suppress the $\overline{Z}_{i}$ error is particularly noteworthy. Error bars are $1\sigma$ standard deviation after bootstrapping the data.
To obtain the fidelity of a logical error, we first observe that each of the logical Bell states is related to the other three logical Bell states through a specific logical error: $\chi^{\prime}=\overline{O}\chi$, where $\overline{O}\in\{I,\overline{X}_{i},\overline{Y}_{i},\overline{Z}_{i}\}$ and $i=1$ or $2$; i.e., a logical operator that acts on either of the two logical qubits. We thus proceed as follows: instead of applying $U_{\rm{enc},\chi}^{\dagger}$, at $t=\tau$ we deliberately unencode into a different logical Bell state $\chi^{\prime}$ [see Fig. 1(a)]. In this manner, each of the even-weight bitstrings gives us a measure of the occurrence of one of the logical errors $\mathcal{E}_{\overline{O},\tau}$, where $\overline{O}$ is determined by the unencoding we choose [see Fig. 1(b,c)]. Using this methodology, we can detect the occurrence of different types of logical errors and quantify the associated error probability.
@@ III Experimental design and methodology
We note that logical state tomography is an alternative method for learning about the performance of LDD, and in particular, for certifying the entanglement of logical Bell states. However, the $[[4,2,2]]$ code imposes some additional challenges in performing logical tomography, which precluded its use in our work; see Appendix D for details.
III.1 Physical dynamical decoupling improves logical Bell state fidelity
We first show that we can substantially improve the logical Bell fidelity by padding the idle gaps of the encoding circuits with physical (as opposed to logical) DD sequences. An idle gap is a temporal circuit segment during which no gates are applied. Such gaps occur, e.g., when a pair of qubits is involved in a two-qubit gate that takes much longer than a single-qubit gate simultaneously being applied to another qubit; the latter is then idle after the completion of the single-qubit gate, while awaiting the completion of the two-qubit gate.
As shown in Fig. 2, we inserted various DD sequences into the idle gaps of Bell state circuits. This includes XY4 [40], universally robust sequences URn [62] for $n=6,8,10,18$, and RGA8a [61]; the former are illustrated along with robust variants in Fig. 1(d-f).
@@ III.1 Physical dynamical decoupling improves logical Bell state fidelity
In all, these results demonstrate that logical Bell state preservation can benefit significantly from physical DD. Having established the utility of physical DD, we proceed to combine it with LDD in the next section.
@@
III.2 Logical error suppression and detection by LDD

### v2

@@ II.4 Illustration of LDD and NDD: the $[[4,2,2]]$ code
The stabilizer group of the $[[4,2,2]]$ code is $\mathcal{S}=\langle XXXX,ZZZZ\rangle$. A canonical set of logical operators for the code can be chosen such that $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, i.e., $\mathcal{L}=\langle i,XIIX,IIXX,IIZZ,ZIIZ\rangle$ (see Appendix B for additional pertinent details regarding this code). Thus, $\{XIIX,IIXX,IIZZ,ZIIZ\}$ is the generator set of the $16$-element LDD group $\tilde{\mathcal{G}}_{\text{LDD}}=\{\tilde{g}_{j}\}$. This LDD group (as well as any of the corresponding NDD groups) suppresses every logical error $E_{\text{log}}$ in the $602^{n-k}(4^{k}-1)=4\times 15=$ 60-element group $(\mathcal{N}(\mathcal{S})\setminus\mathcal{S})/\{\pm 1,\pm i\}$: the group-averaged logical system-bath Hamiltonian $\langle H_{SB}^{\mathcal{L}_{\star}}\rangle_{\mathcal{G}}$ [Eq. 6c] vanishes, as every one of its terms is a logical operator that anticommutes with an element of $\mathcal{G}$.
This leaves $1924^{n}-2^{n+k}=256-64=$ 192 detectable errors. These errors can be written compactly as $\mathcal{D}=\{(XIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\cup(YIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\cup(ZIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\}$, yielding $1923\times|\tilde{\mathcal{N}}(\mathcal{S})|=$ 192 terms. Each such error is detectable because it anticommutes with at least one element of $\mathcal{S}$. Of these, only the $12$ errors of the form $PQPP$, with $P\neq Q\in\{I,X,Y,Z\}$, are unsuppressed by LDD.
Now suppose that instead of suppressing the full $H_{SB}^{\mathcal{L}_{\star}}$, the goal is to suppress just $H_{SB}^{\overline{Z}}\equiv\overline{Z}_{1}\otimes B_{1}^{\text{log}}+\overline{Z}_{2}\otimes B_{2}^{\text{log}}$. This scenario is particularly relevant to the experiments we describe below. In this case, it suffices to use the smaller LDD-subgroup $\langle\overline{X}_{1}\overline{X}_{2}\rangle$, since both $\overline{Z}_{1}$ and $\overline{Z}_{2}$ anticommute with $\overline{XX}\equiv\overline{X}_{1}\overline{X}_{2}$. The corresponding LDD cycle is then $\overline{XX}f_{\tau}\overline{XX}f_{\tau}$, where $f_{\tau}=\exp(-i\tau H_{SB}^{\overline{Z}})$. For crosstalk-related reasons that will become apparent below, it turns out that it is advantageous to replace the second $\overline{XX}$ by $S_{1}\overline{XX}=IXIX\equiv\overline{X^{\prime}X^{\prime}}$, where $S_{1}=XXXX\in\mathcal{S}$. This replacement yields the $4$-pulse NDD cycle $\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{XX}f_{\tau}\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{XX}f_{\tau}=(\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{X^{\prime}X^{\prime}})(S_{1}f_{\tau}S_{1})(\overline{XX}f_{\tau}\overline{XX})(If_{\tau}I)$. This cycle is generated by the NDD group $\tilde{\mathcal{G}}=\langle\overline{XX},S_{1}\rangle=\{I,\overline{XX},S_{1},\overline{X^{\prime}X^{\prime}}\}$. Below, we refer to this decoupling group as NXX.
As another example of NDD, consider the encoded XY4 sequence with the LDD group $\tilde{\mathcal{G}}=\langle\overline{XX},\overline{ZZ}\rangle$ and DD cycle $\overline{YY}f_{\tau}\overline{XX}f_{\tau}\overline{YY}f_{\tau}\overline{XX}f_{\tau}$. Again, for crosstalk-related reasons it will turn out to be advantageous to replace the canonical $\overline{YY}=YIYI$ by $\overline{Y^{\prime}Y^{\prime}}=S_{2}\overline{YY}$, where $S_{2}=YYYY\in\mathcal{S}$. The resulting DD cycle $\overline{Y^{\prime}Y^{\prime}}f_{\tau}\overline{XX}f_{\tau}\overline{Y^{\prime}Y^{\prime}}f_{\tau}\overline{XX}f_{\tau}$ is generated by the NDD group $\tilde{\mathcal{G}}=\langle\overline{XX},\overline{Z^{\prime}Z^{\prime}}\rangle=\{I,\overline{XX},\overline{Y^{\prime}Y^{\prime}},\overline{Z^{\prime}Z^{\prime}}\}$, where $\overline{Z^{\prime}Z^{\prime}}=S_{2}\overline{ZZ}$. Below, we refer to this decoupling group as NXY4.
Figure 1: (a) Schematic of the experimental design: we encode a logical Bell state $\chi$ using the $[[4,2,2]]$ code and let it undergo either free evolution or NDD for a duration $\tau$. We then run the unencoding circuit for $\chi^{\prime}$ and measure the resulting bitstring. The encoding and unencoding circuits together take up to $5\mu$s. (b) The logical error set that relates the four Bell states as a result of a logical operator acting on one ($i=1$ or $2$) of the two logical qubits. (c) Interpretation of different bitstrings for an encoded state ($U_{\rm{enc},\chi}$), depending on the unencoding ($U_{\rm{enc},\chi^{\prime}}^{\dagger}$) circuit used. The color code indicates that the encoded state has undergone either no error (black) or one of the logical errors (red, blue, green) indicated in (b). The DD sequences used in our experiments are (d) Robust Normalizer XX (RNXX), (e) Robust Normalizer XY4 (RNXY4), and (f) physically Staggered XY4 (SXY4). Each panel shows both the physical implementation (top) and the logical interpretation of each sequence (bottom). A dagger in (d) and (e) denotes a negative $\pi$ rotation, i.e., $X=R_{x}(\pi)=\exp(-i\pi\sigma^{x}/2)$, $\tilde{X}=R_{x}(-\pi)=\exp(i\pi\sigma^{x}/2)$, and likewise for $Y$. In (d) the first column is the canonical $\overline{XX}=XIXI$ (see Appendix B), the second column is $\overline{X^{\prime}X^{\prime}}=S_{1}\overline{XX}$, where $S_{1}=XXXX\in\mathcal{S}$, etc. Likewise, in (e), the second column is $\overline{Y^{\prime}Y^{\prime}}=S_{2}\overline{YY}$, where the canonical $\overline{YY}=YIYI$ and $S_{2}=YYYY\in\mathcal{S}$.
III Experimental design
@@
For our experiments, while sequences utilizing the full LDD/NDD group can be constructed (see Appendix C), we instead use a subgroup-NDD, generated by a subgroup of $\mathcal{L}$ and a subgroup of $\mathcal{S}$. This subgroup NDD works better in practice because of its robustness and ability to suppress crosstalk. This section is dedicated to demonstrating experimentally how our hybrid QED + subgroup-NDD protocol results in a significant fidelity enhancement of entangled logical qubits.
The ibm_kyiv processor consists of coupled, fixed-frequency transmons [86]. Such qubits exhibit an always-on interaction between adjacent pairs [87]. This crosstalk gives rise to weight-2 error terms that correspond to logical errors with our $[[4,2,2]]$ code choice. Other errors of weight $\geq 2$, if present, would likewise correspond to logical errors. To demonstrate that NDD can suppress all logical errors, we design an experiment that enables the unequivocal detection of such errors.
To this end, we use the two logical qubits of the $[[4,2,2]]$ code to create logical Bell states |$\ket{\Phi_{\pm}}=(\ket{\overline{00}}\pm\ket{\overline{11}})/\sqrt{2}$ and |$\ket{\Psi_{\pm}}=(\ket{\overline{01}}\pm\ket{\overline{10}})/\sqrt{2}$. The corresponding encoding circuits $U_{\text{enc},\chi}$, where $\chi\in\{\ket{\Phi_{\pm}},\ket{\Psi_{\pm}}\}$, create two copies of physical Bell states, followed by the application of a logical controlled-NOT ($\overline{\text{CNOT}}$) to generate the logical Bell states (see Appendix B). In principle, it would be possible to detect specific logical errors using other logical states (e.g., |$\ket{\overline{00}}$ would allow us to detect $X$- or $Y$-type logical errors), but we specifically opt for logical Bell states as they have shorter encoding circuits. Additionally, the four logical Bell states form an orthonormal basis and the logical errors we are interested in permute this basis.
To estimate the fidelity of a logical Bell state $\chi$, we start from the physical ground state |$\ket{0000}$, encode into the state of interest by applying $U_{\text{enc},\chi}$, and then let it evolve for some time $\tau$, either freely or subject to DD. After time $\tau$ we unencode by applying $U_{\text{enc},\chi}^{\dagger}$, and measure all qubits. Measuring the bitstring $0000$ would indicate that no error occurred.
@@
To obtain the fidelity of a logical error, we first observe that each of the logical Bell states is related to the other three logical Bell states through a specific logical error: $\chi^{\prime}=\overline{O}\chi$, for $\overline{O}\in\{\overline{X}_{i},\overline{Y}_{i},\overline{Z}_{i}\}$ and $i=1$ or $2$; i.e., a logical operator that acts on either of the two logical qubits. More generally, any other combinations of logical weight-$1$ errors also map the input Bell state to the other Bell states. That is, for $\bar{p},\bar{q}\in\{\overline{X},\overline{Y},\overline{Z}\}$, we have $\chi^{\prime}=(\bar{p}\otimes\bar{q})\chi=I\otimes(\bar{p}\bar{q})\chi=(\bar{p}\bar{q})\otimes I\chi$, up to overall phases. We note that the following two-qubit logical errors leave the logical Bell states unchanged and therefore do not influence the experimental outcome: $\overline{O}\chi=\chi$, for $\overline{O}\in\{II,\overline{X}\overline{X},\overline{Y}\overline{Y},\overline{Z}\overline{Z}\}$.
We thus proceed as follows: instead of applying $U_{\rm{enc},\chi}^{\dagger}$, at $t=\tau$ we deliberately unencode into a different logical Bell state $\chi^{\prime}$ [see Fig. 1(a)]. In this manner, each of the even-weight bitstrings gives us a measure of the occurrence of one of the logical errors $\mathcal{E}_{\overline{O},\tau}$, where $\overline{O}$ is determined by the unencoding we choose [see Fig. 1(b,c)]. Using this methodology, we can detect the occurrence of different types of logical errors and quantify the associated error probability.
@@
We note that logical state tomography is an alternative method for learning about the performance of LDD or NDD, and in particular, for certifying the entanglement of logical Bell states. However, the $[[4,2,2]]$ code imposes some additional challenges in performing logical tomography, which precluded its use in our work; see Appendix D for details.
Figure 2: Circuit schematic for encoding |$\ket{\Phi_{+}}$ and unencoding |$\ket{\Phi_{-}}$ with physical and logical DD. Physical XY4 dynamical decoupling sequences (yellow boxes) are inserted into the idle gaps of the circuit. The code subspace is protected with NDD sequences. This is the DD-protected version of the circuit shown in Fig. 1(a).
Figure 3: Performance of NDD (dataset 1). The experiments encode the logical Bell states |$\ket{\Phi_{+}}$ in (a) and (b), and |$\ket{\Psi_{+}}$ in (c) and (d). We then unencode in the logical Bell state related through $\overline{Z}_{i}$ on either of the logical qubits; i.e., |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ in (a) and (b), and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ in (c) and (d). In this setting, the frequency of $0110$ outcomes [(a) and (c)] is a measure of the fidelity of the originally encoded state, while $0000$ corresponds to the detection of a $\overline{Z}_{i}$ error [(b) and (d)]. Without any DD, the fidelity is low [green in (a) and (c)] and the probability of logical errors is high [green in (b) and (d)]. Standard NDD sequences (NXX and NXY4; dashed light red and blue) improve the fidelity and logical error probability at short times ($\lesssim 15\mu$s) but their performance declines at longer times: they exhibit coherent errors as seen in the oscillations of the corresponding curves. Robust sequences (RNXX and RNXY4; solid dark red and blue) are the top performers; their ability to suppress the $\overline{Z}_{i}$ error is particularly noteworthy. Error bars are $1\sigma$ standard deviation after bootstrapping the data.
IV Experimental Results
IV.1 Physical dynamical decoupling improves logical Bell state fidelity
We first show that we can substantially improve the logical Bell fidelity by padding the idle gaps of the encoding circuits with physical (as opposed to logical or normalizer) DD sequences. An idle gap is a temporal circuit segment during which no gates are applied. Such gaps occur, e.g., when a pair of qubits is involved in a two-qubit gate that takes much longer than a single-qubit gate simultaneously being applied to another qubit; the latter is then idle after the completion of the single-qubit gate, while awaiting the completion of the two-qubit gate.
As shown in Fig. 2, we inserted various DD sequences into the idle gaps of Bell state circuits. This includes XY4 [40], universally robust sequences URn [62] for $n=6,8,10,18$, and RGA8a [61].
Since the URn sequences rely heavily on the use of both positive and negative rotations, it is important to clarify that henceforth, we use the notation $X\equiv R_{x}(\pi)=\exp(-i\pi\sigma^{x}/2)$, $\tilde{X}\equiv R_{x}(-\pi)=\exp(i\pi\sigma^{x}/2)$, and likewise for $Y$, to denote physical DD pulses. That is, unlike Section II, the symbols $X$ and $Y$ no longer denote the Pauli $\sigma^{x}$ and $\sigma^{y}$ operators. The reason this matters is that physical DD pulses suffer from control errors (both axis-angle and angle-magnitude errors), as well as errors due to the presence of the system-bath interaction while the pulse is on. In other words, the notation $R_{x}(\pi)$, etc., hides the fact that in reality the exact rotation axis deviates from $x$ and the exact rotation angle deviates from $\pi$. When these deviations are accounted for, DD sequences that are formally identical due to phase cancellation, such as $XX$ and $X\tilde{X}$, result in different operations at the physical level. For more details, see Appendix E.
@@ IV.1 Physical dynamical decoupling improves logical Bell state fidelity
In all, these results demonstrate that logical Bell state preservation can benefit significantly from physical DD. Having established the utility of physical DD, we proceed to combine it with NDD in the next section.
@@
IV.2 Logical error suppression and detection by NDD